# Learning to Segment for Vehicle Routing Problems

### v1

@@ A.2 Proof of FSTA
A.2.1 CVRP
Aggregation Strategy (Two Hypernodes). The detailed implementation of FSTA on CVRP can be found in Appendix A.1.5 and Table 6. Notice that one single hypernode aggregation is also applicatable for CVRP, and $\tilde{d}_{j}$,$\tilde{d_{k}}$ could take other values as long as $\tilde{d}_{j}+\tilde{d_{k}}=d_{j}+...+d_{k}$.
@@ A.2.1 CVRP
A.2.2 VRPTW
Aggregation Strategy (Mixed Strategies). The detailed implementation of FSTA on VRPTW can be found in Appendix A.1.5 and Table 6. We denote $s^{*}_{m}=s_{m}+\text{dist}(x_{m},x_{m+1})$ for $j\leq m<k$ and $s^{*}_{k}=s_{k}$. We further set the service time by $\tilde{s}_{m}=\sum\limits_{m\leq q\leq k}s^{*}_{q}$, and we repeat the temporal time window $[\bar{t}^{l}_{j},\bar{t}^{r}_{j}]$ (which could be infeasible) defined by the following recursive relationship:
$\bar{t}_{m}^{l}$ =$=\begin{cases}t_{k}^{l}&m=k\\ \max\{t_{m}^{l},\bar{t}_{m+1}^{l}-s^{*}_{m}\}&j\leq m\leq k-1,\end{cases}$ (8)
$\bar{t}_{m}^{r}$ =$=\begin{cases}t_{k}^{r}&m=k\\ \min\{t_{m}^{r},\bar{t}_{m+1}^{r}-s^{*}_{m}\}&j\leq m\leq k-1,\end{cases}$
where $[t_{m}^{l},t_{m}^{r}]$ is the time window for a node $x_{m}$, $s_{m}$ is the service time at node $x_{m}$ and $\text{dist}(x_{m},x_{m+1})$ is the time to travel from node $x_{m}$ to node $x_{m+1}$.
Feasibility Proof [Time Window Constraint]. We first prove for the condition that the temporal time window $[\bar{t}^{l}_{j},\bar{t}^{r}_{j}]$ is feasible ($\bar{t}^{l}_{j}<\bar{t}^{r}_{j}$) and single hypernode aggregation is applied. Then, we extend to the infeasible temporal time window condition where dual hypernode aggregation is applied.
@@ A.2.2 VRPTW
  • Inductive Step (segment length = $(k-m)+1>1$). textit the aggregation of the segment $S_{m+1,k}=\left(x_{m+1}\rightarrow...\rightarrow x_{k}\right)$ into $\tilde{S}_{m+1,k}=\{\tilde{x}_{m+1,k}\}$ satisfies condition (1). We want to show that the aggregation of the segment $S_{m,k}=\left(x_{m}\rightarrow...\rightarrow x_{k}\right)$ into $\tilde{S}_{m,k}=\{\tilde{x}_{m,k}\}$ also satisfies condition (1).
@@
Proof of Condition (2): For all $m$, suppose we arrive at the hypernode $\tilde{x}_{m,k}$ at time $t\leq\bar{t}_{m}^{r}$ in the aggregated solution. By definition, the total time spent on the aggregated segment (sum of the waiting time, service time, and the travel time) can be written as the following linear function with $-1$ slope as shown in the first figure in Figure 9.
@@
  • Inductive Step (segment length = $(k-m)+1>1$). Again, suppose the total time spent on the segment $S_{m+1,k}=\left(x_{m+1}\rightarrow...\rightarrow x_{k}\right)$ into $\tilde{S}_{m+1,k}=\{\tilde{x}_{m+1,k}\}$ satisfies the function
@@
      $g_{m}(t)$ =$=\begin{cases}s^{*}_{m}+\tilde{s}_{m+1}&t\geq\bar{t}_{m}^{l}\\ t_{m}^{l}-t+s^{*}_{m}+\tilde{s}_{m+1}&t<\bar{t}_{m}^{l}.\end{cases}$ (13)
      =$=\begin{cases}\tilde{s}_{m}&t\geq\bar{t}_{m}^{l}\\ t_{m}^{l}-t+\tilde{s}_{m}&t<\bar{t}_{m}^{l}\end{cases}=\tilde{g}_{m}(t).$
@@
      =$=\begin{cases}\tilde{s}_{m+1},&t+s^{*}_{m}\geq\bar{t}_{m+1}^{l}\\ \bar{t}_{m+1}^{l}-(t+s^{*}_{m})+\tilde{s}_{m+1},&t+s^{*}_{m}<\bar{t}_{m+1}^{l}\end{cases}$
      =$=\begin{cases}\tilde{s}_{m+1},&t\geq\bar{t}_{m}^{l}\\ \bar{t}_{m}^{l}-t+\tilde{s}_{m+1},&t_{m}^{l}\leq t<\bar{t}_{m}^{l}\end{cases}$
@@
      $g_{m}(t)$ =$=\begin{cases}s^{*}_{m}+\tilde{s}_{m+1}&t\geq\bar{t}_{m}^{l}\\ s^{*}_{m}+\bar{t}_{m}-t+\tilde{s}_{m+1}&t_{m}^{l}\leq t<\bar{t}_{m}^{l}\\ t_{m}^{l}-t+s^{*}_{m}+\bar{t}_{m}^{l}-t_{m}^{l}+\tilde{s}_{m+1}&t<t_{m}^{l},\end{cases}$ (16)
      =$=\begin{cases}\tilde{s}_{m}&t\geq t_{m}^{l}\\ \bar{t}_{m}-t+\tilde{s}_{m}&t_{m}^{l}\leq t<\bar{t}_{m}^{l}\\ \bar{t}_{m}^{l}-t+\tilde{s}_{m}&t<t_{m}^{l},\end{cases}$
      =$=\begin{cases}\tilde{s}_{m}&t\geq\bar{t}_{m}^{l}\\ \bar{t}_{m}^{l}-t+\tilde{s}_{m}&t<\bar{t}_{m}^{l}\end{cases}=\tilde{g}_{m}(t).$
@@
Monotonicity Proof. For the dual hypernode aggregation, please refer to the Monotonicity Proof in A.2.1. For the single hypernode aggregation, notice that
@@
A.2.3 VRPB
Aggregation Strategy (One Hypernode). The detailed implementation of FSTA on VRPB can be found in Appendix A.1.5 and Table 6.
@@ A.2.3 VRPB
Monotonicity Proof. Please refer to the monotonicity proof of VRPTW in Appendix A.2.2.
Appendix B L2Seg Details
B.1 Input Feature Design Details
@@
Previous works [2, 4, 12] typically utilize only basic input features for routing problems (xy-coordinates and normalized demands for node features, and edge cost for edge features). While neural networks can potentially learn complex patterns from these basic features, tailored feature engineering may lead to enhanced model performance. As illustrated in Appendix A.1, we observe that detecting unstable edges may depend on better capturing local dependencies. We therefore design enhanced node and edge features for our learning task, as shown in Table 7. We also include time windows and service time as node features for VRPTWs.
@@
The percentage of the Knearest nodes 3
@@ B.2 Masking Details
B.3 Training Data Collection Details
@@
Input: Solution distribution $\mathcal{P}$, number of instances $N_{\mathcal{P}}$, backbone solver $BS$, number of iterative steps $T_{IS}$, improvement threshold $\eta_{\text{improv}}$, sample coefficient $\alpha_{AC}$
@@ 19      $E_{\text{unstable}}\leftarrow E_{\text{unstable}}\cup E^{P_{\text{TR}}}_{\text{unstable}}$
Appendix C Experimental and Implementation Details
C.1 Backbone solvers
@@
LKH-3. The Lin-Kernighan-Helsgaun algorithm (LKH-3) [7] represents a strong classical heuristic solver for routing problems, which is widely used in NCO for benchmark. It employs sophisticated $k$-opt moves and effective neighborhood search strategies. For our experiments, we impose time limits rather than local search update limits: 150s and 240s for large-capacity CVRP2k and CVRP5k, respectively, and 2m, 4m, and 10m for VRPTW1k, VRPTW2k, and VRPTW5k, respectively. For small-capacity CVRPs, we adopt the results reported in [26].
@@
LNS. Local Neighborhood Search (LNS) [9] is a powerful decomposition-based metaheuristic that iteratively improves solutions by destructively and constructively exploring defined search neighborhoods. We implement LNS following the approach in [4], where neighborhoods consisting of three adjacent subroutes are randomly selected for re-optimization. We establish time limits of 150s and 240s for large-capacity CVRP2k and CVRP5k, respectively; 2.5m, 4m, and 5m for small-capacity CVRP1k, CVRP2k, and CVRP5k, respectively; and 2m, 4m, and 10m for VRPTW1k, VRPTW2k, and VRPTW5k, respectively. LKH-3 serves as the backbone solver with a 1,000 per-step local search updates limit.
@@
L2D. Learning to Delegate (L2D) [4] is the state-of-the-art learning-based optimization framework that integrates neural networks with classical optimization solvers to intelligently delegate subproblems to appropriate solvers. The framework employs a neural network trained to identify the most promising neighborhoods for improvement. For comparative fairness, we apply identical time limits and backbone solver configurations as used in our LNS implementation. When augmented by L2Seg, training proceeds in two stages: we first train the L2D models following the methodology in [4], then train the L2Seg model using the resulting pre-trained L2D models.
@@
Initial Solution Heuristics. For both training data generation and inference, we employ the initial solution heuristic inspired by [4]. Our method partitions nodes according to their angular coordinates with respect to the depot. We begin by selecting a reference node, marking its angle as 0, and incrementally incorporate additional nodes into the same group until the collective demand approaches the capacity threshold ($\alpha_{\text{init}}K_{\text{veh}}C\approx\sum d_{i}$), where approximately $K_{\text{veh}}$ vehicles would be required to service the group. This process continues sequentially, forming new groups until all customers are assigned. Finally, we apply LKH-3 in parallel to solve each subproblem independently. In our implementation, we set $K_{\text{veh}}=6$ and $\alpha_{\text{init}}=0.95$ as the controlling parameters.
C.2 Baselines
@@
In this section, we provide further clarification regarding the baselines used in our comparative analysis, beyond the backbone solvers. We independently executed LKH-3, LNS, and L2D using consistent parameters. Results for SIL were sourced from [5], L2R from [42], and all other baselines from [26]. When multiple variants of a baseline were presented in the original publications, we selected the configuration that achieved the best objective values.
@@
It is important to note that all reported results were evaluated on identical test instances (for CVRPs) or on instances sampled from the same distribution (for VRPTWs), ensuring fair comparison. Moreover, our experiments were conducted on hardware with less powerful GPUs compared to those utilized in [5, 26, 42]. This hardware discrepancy suggests that the performance advantages demonstrated by our proposed model would likely persist or potentially increase if all methods were evaluated on identical computing infrastructure.
@@
We re-implemented the backbone solvers and L2D [4] to ensure a fair and strong comparison. Notably, prior studies [26, 40] did not explore configurations optimized for L2D’s full potential. Specifically, they imposed overly conservative limits (e.g., only allowing 1 trail) on LKH-3 local search updates and did not supply current solution information to the LKH-3 solver during the resolution process. This significantly weakened L2D’s performance in their benchmarks. In contrast, our comparison reflects L2D’s best achievable performance.
C.3 Instance Generation
In general, we generate all training and test instances following established methodologies: Zheng et al. [26] for CVRP and Solomon [52] for VRPTW. Specifically, For small-capacity CVRPs, nodes are uniformly distributed within the $[0,1]$ square, with integer demands ranging from 1 to 9 (inclusive). Vehicle capacities are set to $C=200$, $300$, and $300$ for problem sizes 1k, 2k, and 5k, respectively. For large-capacity CVRPs, we maintain identical configurations except for increased vehicle capacities of $C=500$ and $1000$ for CVRP1k and CVRP5k, respectively. For VRPTWs, we adopt the same spatial distribution, demand structure, and capacity constraints as the small-capacity CVRPs. Service times are uniformly set to 0.2 time units for each customer and 0 for the depot. Time windows are generated according to the methodology outlined in Solomon [52].
  • Validation: 30 instances per problem configuration
  • Testing: For small-capacity CVRPs, we utilize the 1,000 test instances from Zheng et al. [26]; for large-capacity CVRPs and VRPTWs, we evaluate on 100 instances sampled from the same distribution as the training data
C.4 Parameters and Training Hyperparameters

### v2

@@ B.2 Proof of FSTA
B.2.1 CVRP
Aggregation Strategy (Two Hypernodes). The detailed implementation of FSTA on CVRP can be found in Appendix B.1.5 and Table 6. Notice that one single hypernode aggregation is also applicatable for CVRP, and $\tilde{d}_{j}$$\tilde{d_{k}}$ could take other values as long as $\tilde{d}_{j}+\tilde{d_{k}}=d_{j}+...+d_{k}$.
@@ B.2.1 CVRP
B.2.2 VRPTW
Aggregation Strategy (Mixed Strategies). The detailed implementation of FSTA on VRPTW can be found in Appendix B.1.5 and Table 6. We denote $s^{*}_{m}=s_{m}+\text{dist}(x_{m},x_{m+1})$ for $j\leq m<k$ and $s^{*}_{k}=s_{k}$. We further set the service time by $\tilde{s}_{m}=\sum\limits_{m\leq q\leq k}s^{*}_{q}$, and we repeat the temporal time window $\bar{t}^{l}_{j},\bar{t}^{r}_{j}]$ (which could be infeasible) defined by the following recursive relationship:
[…]
where $t_{m}^{l},t_{m}^{r}]$ is the time window for a node $x_{m}$, $s_{m}$ is the service time at node $x_{m}$ and $\text{dist}(x_{m},x_{m+1})$ is the time to travel from node $x_{m}$ to node $x_{m+1}$.
Feasibility Proof [Time Window Constraint]. We first prove for the condition that the temporal time window $\bar{t}^{l}_{j},\bar{t}^{r}_{j}]$ is feasible ($\bar{t}^{l}_{j}<\bar{t}^{r}_{j}$) and single hypernode aggregation is applied. Then, we extend to the infeasible temporal time window condition where dual hypernode aggregation is applied.
@@ B.2.2 VRPTW
  • Inductive Step (segment length = (km)+1>1(k-m)+1>1). textit the aggregation of the segment $S_{m+1,k}=\left(x_{m+1}\rightarrow...\rightarrow x_{k}\right)$ into $\tilde{S}_{m+1,k}=\{\tilde{x}_{m+1,k}\}$ satisfies condition (1). We want to show that the aggregation of the segment $S_{m,k}=\left(x_{m}\rightarrow...\rightarrow x_{k}\right)$ into $\tilde{S}_{m,k}=\{\tilde{x}_{m,k}\}$ also satisfies condition (1).
@@
Proof of Condition (2): For all $m$, suppose we arrive at the hypernode $\tilde{x}_{m,k}$ at time $t\leq\bar{t}_{m}^{r}$ in the aggregated solution. By definition, the total time spent on the aggregated segment (sum of the waiting time, service time, and the travel time) can be written as the following linear function with 1-1 slope as shown in the first figure in Figure 10.
@@
  • Inductive Step (segment length = (km)+1>1(k-m)+1>1). Again, suppose the total time spent on the segment $S_{m+1,k}=\left(x_{m+1}\rightarrow...\rightarrow x_{k}\right)$ into $\tilde{S}_{m+1,k}=\{\tilde{x}_{m+1,k}\}$ satisfies the function
@@
[…]
@@
[…]
@@
[…]
@@
Monotonicity Proof. For the dual hypernode aggregation, please refer to the Monotonicity Proof in B.2.1. For the single hypernode aggregation, notice that
@@
B.2.3 VRPB
Aggregation Strategy (One Hypernode). The detailed implementation of FSTA on VRPB can be found in Appendix B.1.5 and Table 6.
@@ B.2.3 VRPB
Monotonicity Proof. Please refer to the monotonicity proof of VRPTW in Appendix B.2.2.
B.2.4 1-VRPPD.
Aggregation Strategy (Three Hypernodes). The detailed implementation of FSTA on 1-VRPPD can be found in Appendix B.1.5 and Table 6.
Feasibility Proof [1-Commodity Pickup and Delivery Constraint]. A feasible $\tilde{\mathcal{R}}_{+}$ indicates that whenever the vehicle is traveling an aggregated segment $\tilde{S}_{j,k}$, denoted the starting load of the vehicle to be $d_{\text{st}}$ and ending load of the vehicle to be $d_{\text{ed}}$, we have:
[…]
which requires $D^{\text{min}}\leq d_{\text{st}}\leq C-D^{\text{max}}$ and $d_{\text{ed}}=d_{\text{st}}+D^{k}$.
On the other hand, a feasible solution $\mathcal{R}_{+}$ indicates that whenever the vehicle is traveling a segment $S_{j,k}$, denoted the starting load of the vehicle to be $d_{\text{st}}$ and ending load of the vehicle to be $d_{\text{ed}}$, we have:
$0\leq d_{\text{st}}+D^{i}\leq C,\,\,\,\forall i$ (23)
which also requires $D^{\text{min}}\leq d_{\text{st}}\leq C-D^{\text{max}}$ and $d_{\text{ed}}=d_{\text{st}}+D^{k}$. Then, we have a feasible $\tilde{\mathcal{R}}_{+}\Rightarrow$ a feasible $\mathcal{R}_{+}$.
Monotonicity Proof. As $\text{dist}(\tilde{x}_{j},\tilde{x}_{\text{mid}})=\text{dist}(\tilde{x}_{\text{mid}},\tilde{x}_{k})=0$, we can eliminate the middle hypernode and use a two-hypernode representation when calculating the routing objective. Please refer to the monotonicity proof of CVRP in Appendix B.2.1 for the monotonicity proof of two-hypernode representation.
Appendix C L2Seg Details
C.1 Comparative Analysis of L2Seg Against Existing Methods
Comparisons with Large Neighborhood Search (LNS). (1) LNS (Large Neighborhood Search) operates within a bounded local neighborhood. The algorithm selects a specific region, destroys elements within that boundary, and rebuilds only that portion while keeping the rest of the solution intact. For instance, in Li et al. (2021), LNS selects 3-5 subroutes as its neighborhood, modifying only these routes while leaving all others completely unchanged. There is a clear demarcation between the modified neighborhood and the preserved structure. (2) FSTA (our method), in contrast, operates more globally across the entire solution. It can break existing edges and aggregate segments throughout all subroutes simultaneously, without any predefined neighborhood boundaries. The modifications are distributed across the entire solution rather than confined to a local region, which represents a fundamental departure from existing LNS to more efficiently guide the search. We note that such a flexible framework would not be possible without the proposed ML component, which also constitutes the core novelty and contribution of our work to the field. (3) Moreover, FSTA and LNS are complementary: FSTA can be applied on top of LNS, where LNS first selects a large neighborhood, then FSTA fixes stable edges globally within that selected region.
Comparisons with Evolutionary Algorithms. L2Seg framework and evolutionary algorithms (Vidal, 2022)) approach the preservation of solution components from different angles and with distinct goals, and are not interchangeable in use. Evolutionary algorithms (Vidal, 2022)) rely on crossover to merge relatively “good” components from different parents, aiming to promote diversity and generate promising offspring, while our L2Seg framework introduces a learning-guided mechanism to detect unstable edges and aggregates stable edge sequences into hypernodes, enabling a new form of segment-based decomposition that improves scalability and efficiency.
Comparisons with Path Decomposition Method. (1) Firstly, path decomposition relies on geometric heuristics (e.g., clustering routes by barycenter distances) to identify decomposition boundaries. In contrast, L2Seg employs deep learning models (synergistic NAR-AR architecture) to intelligently predict which segments should be aggregated, capturing complex patterns that simple heuristics cannot identify. We also propose a novel learning-guided framework with bespoke training and inference processes that are unique to the machine learning method. (2) Secondly, while some prior work explores similar decomposition ideas (e.g., on CVRP only), we are the first to study FSTA decomposition theoretically, providing formal definitions, feasibility theorems, and monotonicity guarantees for various VRPs. (3) Lastly, we empirically demonstrate that by leveraging deep learning in our L2Seg framework, our method consistently achieves significant speedups on state-of-the-art backbones. This provides new insights for the community, highlighting the power of learning-guided optimization in accelerating combinatorial solvers.
Comparisons with Previous Learning-based Framework L2D (Li et al., 2021). (1) Different from the sub-route level, our method detects unstable edges both within and across sub-routes, enabling more global and flexible decomposition. (2) It optimizes beyond localized neighborhoods by identifying improvements that span multiple distant regions simultaneously. (3) It reduces the size of sub-routes by aggregating stable segments into hypernodes, whereas L2D reduces only the number of sub-routes per iteration. This segment-level aggregation allows more adaptive and coarse-grained reduction, offering higher efficiency and solution quality, while remaining complementary to L2D.
C.2 Input Feature Design Details
@@
Previous works Kool et al. (2018), Li et al. (2021), Kwon et al. (2020) typically utilize only basic input features for routing problems (xy-coordinates and normalized demands for node features, and edge cost for edge features). While neural networks can potentially learn complex patterns from these basic features, tailored feature engineering may lead to enhanced model performance. As illustrated in Appendix B.1, we observe that detecting unstable edges may depend on better capturing local dependencies. We therefore design enhanced node and edge features for our learning task, as shown in Table 7. We also include time windows and service time as node features for VRPTWs.
@@
The percentage of the K% nearest nodes 3
@@ C.3 Masking Details
C.4 Training Data Collection Details
@@
Input: Solution distribution $\mathcal{P}$, number of instances $N_{\mathcal{P}}$, backbone solver BSBS, number of iterative steps $T_{IS}$, improvement threshold $\eta_{\text{improv}}$, sample coefficient $\alpha_{AC}$
@@ 19 $E_{\text{unstable}}\leftarrow E_{\text{unstable}}\cup E^{P_{\text{TR}}}_{\text{unstable}}$
Appendix D Experimental and Implementation Details
D.1 Backbone solvers
@@
LKH-3. The Lin-Kernighan-Helsgaun algorithm (LKH-3) Helsgaun (2017) represents a strong classical heuristic solver for routing problems, which is widely used in NCO for benchmark. It employs sophisticated $k$-opt moves and effective neighborhood search strategies. For our experiments, we impose time limits rather than local search update limits: 150s and 240s for large-capacity CVRP2k and CVRP5k, respectively, and 2m, 4m, and 10m for VRPTW1k, VRPTW2k, and VRPTW5k, respectively. For small-capacity CVRPs, we adopt the results reported in Zheng et al. (2024).
@@
LNS. Local Neighborhood Search (LNS) Shaw (1998) is a powerful decomposition-based metaheuristic that iteratively improves solutions by destructively and constructively exploring defined search neighborhoods. We implement LNS following the approach in Li et al. (2021), where neighborhoods consisting of three adjacent subroutes are randomly selected for re-optimization. We establish time limits of 150s and 240s for large-capacity CVRP2k and CVRP5k, respectively; 2.5m, 4m, and 5m for small-capacity CVRP1k, CVRP2k, and CVRP5k, respectively; and 2m, 4m, and 10m for VRPTW1k, VRPTW2k, and VRPTW5k, respectively. LKH-3 serves as the backbone solver with a 1,000 per-step local search updates limit.
@@
L2D. Learning to Delegate (L2D) Li et al. (2021) is the state-of-the-art learning-based optimization framework that integrates neural networks with classical optimization solvers to intelligently delegate subproblems to appropriate solvers. The framework employs a neural network trained to identify the most promising neighborhoods for improvement. For comparative fairness, we apply identical time limits and backbone solver configurations as used in our LNS implementation. When augmented by L2Seg, training proceeds in two stages: we first train the L2D models following the methodology in Li et al. (2021), then train the L2Seg model using the resulting pre-trained L2D models.
@@
Initial Solution Heuristics. For both training data generation and inference, we employ the initial solution heuristic inspired by (Li et al., 2021). Our method partitions nodes according to their angular coordinates with respect to the depot. We begin by selecting a reference node, marking its angle as 0, and incrementally incorporate additional nodes into the same group until the collective demand approaches the capacity threshold ($\alpha_{\text{init}}K_{\text{veh}}C\approx\sum d_{i}$), where approximately $K_{\text{veh}}$ vehicles would be required to service the group. This process continues sequentially, forming new groups until all customers are assigned. Finally, we apply LKH-3 in parallel to solve each subproblem independently. In our implementation, we set $K_{\text{veh}}=6$ and $\alpha_{\text{init}}=0.95$ as the controlling parameters.
D.2 Baselines
@@
In this section, we provide further clarification regarding the baselines used in our comparative analysis, beyond the backbone solvers. We independently executed LKH-3, LNS, and L2D using consistent parameters. Results for SIL were sourced from Luo et al. (2024), L2R from Zhou et al. (2025a), and all other baselines from Zheng et al. (2024). When multiple variants of a baseline were presented in the original publications, we selected the configuration that achieved the best objective values. Since the original implementation of NDS (Hottung et al., 2025) was evaluated on NVIDIA A100 GPUs whereas our experiments use NVIDIA V100 GPUs, we re-ran NDS on our hardware for fair comparison.
@@
It is important to note that all reported results were evaluated on identical test instances (for CVRPs) or on instances sampled from the same distribution (for VRPTWs), ensuring fair comparison. Moreover, our experiments were conducted on hardware with less powerful GPUs compared to those utilized in Luo et al. (2024), Zheng et al. (2024), Zhou et al. (2025a). This hardware discrepancy suggests that the performance advantages demonstrated by our proposed model would likely persist or potentially increase if all methods were evaluated on identical computing infrastructure.
@@
We re-implemented the backbone solvers and L2D (Li et al., 2021) to ensure a fair and strong comparison. Notably, prior studies (Zheng et al., 2024, Ye et al., 2024) did not explore configurations optimized for L2D’s full potential. Specifically, they imposed overly conservative limits (e.g., only allowing 1 trail) on LKH-3 local search updates and did not supply current solution information to the LKH-3 solver during the resolution process. This significantly weakened L2D’s performance in their benchmarks. In contrast, our comparison reflects L2D’s best achievable performance.
D.3 Parameters and Training Hyperparameters